1a. Suppose you are interested in investing in a tax-free municipal bond having a maturity value of $1,000 in nine months (or exactly 273.75 days which is 0.75 of one year). This bond does not pay any interest (or coupon) payments but rather sells at a discount price of $900. What is your yield to maturity over this 9-month holding period (iYTM)?

1b. Now suppose you are considering a U.S. Treasury bond of identical risk and time to maturity, however the T-bond is a taxable security. Based on your YTM from the municipal bond above and the fact that you are in a 22% marginal tax bracket, what must be the pre-tax YTM on the T-bond in order to make you indifferent between the two bonds?
Answer #1

a)

Interest earned in 0.75 year=1000-900=$100

Principal investment=P=$900

Time period=t=0.75 year

YTM=r=?

We know that

Simple interest=P*r*t

100=900*r*0.75

r=100/(900*0.75)=14.8148%

b)

Pretax YTM=r/(1-Tax rate)=14.8148%/(1-22%)=18.9934%